3. Gains from trade Consider two neighboring island countries called Euphoria and Contente. They each have 4 million labor hoIn the following table, enter each countrys production decision on the third row of the table (marked Production). Suppose
Answer #1

Blanks-

1) 16/4 = 4

2) 12/6 = 2

3) Contente

4) Euphoria

5) 6*4 = 24

6) 16*4 = 64

7) 24-18 = 6

8) 64-52 = 12

Euphoria

Contente

Jeans

Corn

Jeans

Corn

(Millions of pairs)

(Millions of bushels)

(Millions of pairs)

(Millions of bushels)

Without Trade

Production

12

16

6

36

Consumption

12

16

6

36

With Trade

Production

0

64

24

0

Trade action

Import 14

Export 42

Export 14

Import 42

Consumption

14

22

10

42

Gains from Trade

Increase in Consumption

2

6

4

6